Carol, a single TP, earns $75,000 in taxable income and $10,000 in interest from an investment in the City of Chicago bonds (non-taxable). Using the U.S. tax rate schedule:

  1. How much federal income tax will she owe?
  2. What is her average tax rate?
  3. What is her effective tax rate?
  4. If Carol earns an additional $40,000 of taxable income, what is her marginal tax rate?
  5. What if she had $40,000 of additional deductions instead?
Answer #1
In the Books of Carol Co.
= 4991.25 + 25% 38750 14678.75 Fedral Income
= In this Problem Carol Fit into 25% Capacity U.S. Tax Rate Schedule
Particulars Formula Solutions Percentage
Avg Tax Rate = Total Tax = 14678.75 * 100 19.57167 %
Taxable Income 75000
Effective tax rate = Total Tax = 14678.75 * 100 17.26912 %
Taxable Income (75000+10000)
A. Marginal Tax Rate = Total Tax = (25493.25-14678.75) * 100 27.03625 %
Taxable Income (115000-75000)
B. Marginal Tax Rate = Total Tax = (4803.75-14678.75) * 100 24.6875 %
Taxable Income (35000-75000)
Note : All amounts are in $